# Dissipative solutions to Stochastic 3D Euler equations

### v2

@@ 3. Construction of the velocity perturbation
This section explains the construction of the perturbation $w:=v_{q+1}-v_{q}$, where $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$ is a solution to (2.13). In particular, $w$ is obtained by adapting the construction introduced in [DLK23] to our stochastic framework.
@@ 3.1. Mikado flows
Suppose that $\mathcal{F_{\phi}}:=\{k_{1},k_{2},k_{3}\}\subset\mathbb{Z}^{3}$ is an orthogonal frame and define
(3.3) $k_{4}:=-\sum\limits_{i=1}^{3}k_{i}.$
Then, for any $N_{0}>0$, there exist real-valued affine functions $\{\Lambda_{k_{i}}\}_{i=1}^{4}$ on $\mathbb{R}^{3}$ such that their restrictions to $\mathcal{V}_{N_{0}}:=\{u\in\mathbb{R}^{3}:|u|\leq N_{0}\}$ are of class $\mathscr{C}^{\infty}(\mathcal{V}_{N_{0}},[N_{0},\infty))$ and for every $u\in\mathcal{V}_{N_{0}}$ it holds
@@
As in [DLK23], these two lemmas allow to find 27 disjoint families of directions $\{\mathcal{F}^{j}\}_{j\in\mathbb{Z}^{3}_{3}}$, indexed by equivalence classes $j\in\mathbb{Z}^{3}_{3}$ (triples of integers modulo $3$), such that any $\mathcal{F}^{j}$ can be decomposed as $\mathcal{F}^{j}=\mathcal{F}^{j}_{R}\cup\mathcal{F}^{j}_{\phi}$, where $\mathcal{F}^{j}_{R}$ and $\mathcal{F}^{j}_{\phi}$ satisfy the assumptions of Lemmas˜3.1 and 3.2, respectively. For any direction $f\in\mathcal{F}:=\bigcup_{j\in\mathbb{Z}^{3}_{3}}\mathcal{F}^{j}$ let us define
@@
which is nothing but the the line in direction $f$ passing through the origin, but seen as subset of $\mathbb{T}^{3}$. Moreover, le $\eta$ be a positive constant and $\psi_{f}\in C^{\infty}(\mathbb{T}^{3},\mathbb{R})$ be a function such that:
@@ 3.2. Partition of unity and shifts
Next, let $\nu:=\frac{1-7\alpha}{3\alpha}$ such that $(7+\nu)\alpha<1$ and $(4+2\nu)\alpha-1<-3\alpha$. Then, we introduce the parameters
(3.10) $\epsilon:=\begin{cases}C_{\epsilon}\lambda_{0}^{h_{5}}\lambda_{1}^{h_{6}},\,&q=0,\\ C_{\epsilon}\lambda_{q}^{h_{1}}\lambda_{q+1}^{h_{2}}\delta_{q}^{h_{3}}\delta_{q+1}^{h_{4}},\,&q\geq 1,\end{cases}\qquad C_{\epsilon}:=(C_{v}\overline{M}+L)^{-1}\min\left(\frac{1}{2},\frac{\inf_{j}N_{0}(\mathcal{F}_{R}^{j})}{6}\right),$
which represents the size of the temporale partition, and
(3.11) $\mu:=\begin{cases}\lambda_{0}^{u_{5}}\lambda_{1}^{u_{6}},\,&q=0\\ \lambda_{q}^{u_{1}}\lambda_{q+1}^{u_{2}}\delta_{q}^{u_{3}}\delta_{q+1}^{u_{4}},\,&q\geq 1\end{cases},$
which represents instead the size of the spatial partition, where the exponents $h_{i},u_{i}$ are given by:
(3.12) $h_{1}=u_{1}=-1,\quad h_{2}=u_{2}=0,\quad h_{3}=u_{3}=-(2+\nu),\quad u_{4}=-u_{3},\quad h_{4}=\frac{3+2\nu}{2},$
$h_{5}=u_{5}=-1/2,\quad h_{6}=u_{6}=-1/2.$
Indeed, given index sets
@@
(3.15) $\mathcal{F}[\mathbf{P}_{\leq 2^{j}}f](\xi):=\mathfrak{m}(\xi\,2^{-j})\mathcal{F}[f](\xi),\quad\mathcal{F}[\mathbf{P}_{>2^{j}}f](\xi):=[1-\mathfrak{m}(\xi\,2^{-j})]\mathcal{F}[f](\xi),$
The symbol $\mathcal{F}$ here indicated the Fourier transform and should not be confused with the set of directions introduced in Section˜3.1. Given the vector $(e_{1},e_{2},e_{3},e_{4}):=(-1,0,-\frac{4+\nu}{2},\frac{4+\nu}{2}),(e_{5},e_{6}):=(-1/2,-1/2)$, let us introduce the frequency cutoff parameter
(3.16) $l:=\begin{cases}\lambda_{0}^{e_{5}}\lambda_{1}^{e_{6}},\,&q=0\\ \lambda_{q}^{e_{1}}\lambda_{q+1}^{e_{2}}\delta_{q}^{e_{3}}\delta_{q+1}^{e_{4}},\quad&q\geq 1\\ \end{cases},$
@@
(3.18) $\mathbf{P}_{2^{j}}f:=\mathbf{P}_{>2^{j-1}}-\mathbf{P}_{>2^{j}},$
that will be useful later. Then we define the spatial regularization of our velocity fields in the following way:
@@ 3.3. Perturbation of the velocity
where the index $o$ stands for oscillatory. The amplitudes $a_{I}$ will be defined in the following Section˜3.4, and $\psi_{I}$ is given by
@@
The proof of the existence of such a family of shifts is a simple adaptation of [DLK23, Proposition 3.5], with only minor modifications needed to take the randomness of $u_{l}$ into account, and is omitted for the sake of brevity. As in [DLK23], we can expand $\psi_{I},\psi^{2}_{I},\psi^{3}_{I}$ in Fourier series as follows:
(3.23) $\psi_{I}=\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{b}_{I,k}e^{ik\cdot x},\quad\psi^{2}_{I}=\mathring{c}_{I,0}+\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{c}_{I,k}e^{ik\cdot x},\quad\psi^{3}_{I}=\mathring{d}_{I,0}+\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{d}_{I,k}e^{ik\cdot x},$
@@
(3.24) $w_{o}$ $=\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}b_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
(3.25) $w_{o}\otimes w_{o}$ $=c_{0}+\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}c_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
(3.26) $\frac{|w_{o}|^{2}}{2}w_{o}$ $=d_{0}+\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}d_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
@@
(3.27) $b_{m,k}=\theta_{m}(t)\sum_{I\in\mathcal{I}_{m}}\chi_{I}(\xi_{I}(t,x))a_{I}(x,t)\overline{f_{I}}(x,t)\,\mathring{b}_{I,k},$
(3.28) $c_{m,k}=\theta^{2}_{m}(t)\sum_{I\in\mathcal{I}_{m}}\chi^{2}_{I}(\xi_{I}(x,t))a_{I}^{2}(t,x)\left(\overline{f_{I}}\otimes\overline{f_{I}}\right)(t,x)\,\mathring{c}_{m,k},$
(3.29) $d_{m,k}=\theta^{3}_{m}(t)\sum_{I\in\mathcal{I}_{m}}\chi^{3}_{I}(\xi_{I}(x,t))a_{I}^{3}(t,x)\left(\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}\right)(t,x)\,\mathring{d}_{m,k},$
@@
(3.31) $c_{0}=\sum_{I\in\mathcal{I}}\theta^{2}_{I}(t)\chi_{I}^{2}(\xi(x,t))a_{I}^{2}(t,x)\left(\overline{f_{I}}\otimes\overline{f_{I}}\right)(t,x)\fint_{\mathbb{T}^{3}}\psi_{I}(t,y)^{2}dy,$
(3.32) $d_{0}=\sum_{I\in\mathcal{I}}\theta^{3}_{I}(t)\chi_{I}^{3}(\xi(x,t))a_{I}^{3}(t,x)\left(\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}\right)(t,x)\fint_{\mathbb{T}^{3}}\psi_{I}(t,y)^{3}dy.$
@@
(3.33) $w_{c}:=$ $\frac{1}{\lambda_{q+1}}\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}e_{m,k}e^{i\lambda_{q+1}\xi_{m}(x,t)\cdot k},$
(3.34) $e_{m,k}=$ $\sum_{I\in\mathcal{I}_{m}}\nabla\left(\theta_{m}(t)\chi_{I}(\xi_{I}(t,x))a_{I}(t,x)\mathring{b}_{I,k}\right)\times\left(\nabla\xi_{I}^{T}(t,x)\frac{ik\times f_{I}}{|k|^{2}}\right).$
@@ 3.4. Choice of the amplitudes
(3.35) $l_{\rm temp}:=\begin{cases}\lambda_{0}^{s_{5}}\lambda_{1}^{s_{6}},\,&q=0\\ \lambda_{q}^{s_{1}}\lambda_{q+1}^{s_{2}}\delta_{q}^{s_{3}}\delta_{q+1}^{s_{4}},\,&q\geq 1\end{cases},$
with $(s_{1},s_{2},s_{3},s_{4})=(-1,0,-\frac{4+\nu}{2},\frac{3+\nu}{2}),\,(s_{5},s_{6})=(-1/2,-1/2)$. The asymmetric choice for the support of $\eta$ guarantees that convolutions with $\eta$ preserve $\{\mathbb{F}_{t}\}_{t\geq 0}$ adaptedness. Then, define
(3.36) $R_{l}(t,x)$ $=\int_{\mathbb{R}}(\mathbf{P}_{\leq l^{-1}}R_{q})\left(t+s,\mu(t;t+s,x)\right)\eta_{l_{\rm temp}}(s)ds,$
(3.37) $\phi_{l}(t,x)$ $=\int_{\mathbb{R}}(\mathbf{P}_{\leq l^{-1}}\phi_{q})\left(t+s,\mu(t;t+s,x)\right)\eta_{l_{\rm temp}}(s)ds,$
@@
(3.40) $a_{I}(t,x):=|\overline{f_{I}}|^{-2/3}2^{1/3}\Lambda_{f_{I}}^{1/3}\left(-\frac{\nabla\xi_{I}(t,x)\phi_{l}(t,x)}{\overline{M}^{3}\lambda_{q}^{-3\gamma/2}\delta_{q+1}^{3/2}}\right)\overline{M}(\lambda_{q}^{-3\gamma/2}\delta_{q+1}^{3/2})^{1/3}.$
Notice that $\epsilon\|u_{l}(t)\|_{1}<1/2$ implies $|\overline{f}_{I}|>1/2$ by (3.30). Then, (3.8) and (3.32) plus the disjoint supports property gives
(3.41) $d_{0}$ $=\sum_{I\in\mathcal{I_{\phi}}}\theta^{3}_{I}(t)\chi_{I}^{3}(\xi(x,t))a_{I}^{3}(t,x)\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}(t,x)$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))\nabla\xi_{m}^{-1}(t,x)\sum_{\begin{subarray}{c}I\in\mathcal{I}_{\phi}\\ I=(m,n,f_{I})\end{subarray}}\overline{M}^{3}\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}\Lambda_{f_{I}}\left(-\frac{\nabla\xi_{I}(t,x)\phi_{l}(t,x)}{\overline{M}^{3}\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}}\right)f_{I}$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))(-\phi_{l}(t,x))=-\phi_{l}(t,x),$
where the second-to-last equality is due to Lemma˜3.2 applied with $N_{0}=e$, while the last one is due to the space and time partitions of unity.
@@
(3.42) $a_{I}(t,x)$ $:=\Gamma_{f}\left(\mathrm{Id}+(\nabla\xi_{I}\nabla\xi_{I}^{T}-\mathrm{Id})-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\rho^{1/2},$
@@
where $\Gamma_{f}$ and $N_{0}(\mathcal{F}_{R}^{j})$ are given by Lemma˜3.1, $C>0$ is a constant depending on the $\Gamma_{f}$’s functions and $\sup_{f\in\mathcal{F}}|f|_{\infty}$, and
(3.44) $M_{m,n}:=\sum_{\begin{subarray}{c}J\in I_{\phi},\\ J=(m,n,f_{J}),\\ |m-m^{\prime}|_{\infty}\leq 1,\\ |n-n^{\prime}|_{\infty}\leq 1\end{subarray}}a_{J}^{2}(t,x)\chi_{J}^{2}(\xi_{J}(t,x))\theta_{J}^{2}(t)\left(\overline{f_{J}}\otimes\overline{f_{J}}\right)(t,x).$
@@
(3.45) $c_{0}$ $=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))\left[\sum_{I\in\mathcal{I}_{m,n,R}}a_{I}^{2}(t,x)\nabla\xi_{I}^{-1}(t,x)(f_{I}\otimes f_{I})\nabla\xi_{I}^{-T}(t,x)+M_{m,n}\right]$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))\left[\rho\nabla\xi_{I}^{-1}\left(\nabla\xi_{I}\nabla\xi_{I}^{T}-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\nabla\xi_{I}^{-T}+M_{m,n}\right]$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t)\chi_{n}^{6}(\xi_{m}(x,t))\left[\rho\mathrm{Id}-R_{l}\right]=\rho\mathrm{Id}-R_{l},$
where the second to last equality comes from Lemma˜3.1 under the additional assumption

### v3

@@ 3. Construction of the velocity perturbation
This section explains the construction of the perturbation $w:=v_{q+1}-v_{q}$, where $(v_{q},p_{q},R_{q},\phi_{q},z_{q})$ is a solution to (2.9). In particular, $w$ is obtained by adapting the construction introduced in [DLK23] to our stochastic framework, to which we refer for more details.
@@ 3.1. Mikado flows
Let $\mathcal{F_{\phi}}:=\{k_{i}\}_{i=1}^{4}\subset\mathbb{Z}^{3}$ be a family of vectors such that $\{k_{1},k_{2},k_{3}\}\subset\mathbb{Z}^{3}$ is an orthogonal frame and $k_{4}=-(k_{1}+k_{2}+k_{3})$. Then, for any $N_{0}>0$, there exist real-valued affine functions $\{\Lambda_{k_{i}}\}_{i=1}^{4}$ on $\mathbb{R}^{3}$ such that their restrictions to $\mathcal{V}_{N_{0}}:=\{u\in\mathbb{R}^{3}:|u|\leq N_{0}\}$ are of class $\mathscr{C}^{\infty}(\mathcal{V}_{N_{0}},[N_{0},\infty))$ and for every $u\in\mathcal{V}_{N_{0}}$ it holds
@@
As in [DLK23], these two lemmas allow to find 27 disjoint families of directions $\{\mathcal{F}^{j}\}_{j\in\mathbb{Z}^{3}_{3}}$, indexed by equivalence classes $j\in\mathbb{Z}^{3}_{3}$ (triples of integers modulo $3$), such that any $\mathcal{F}^{j}$ can be decomposed as $\mathcal{F}^{j}=\mathcal{F}^{j}_{R}\cup\mathcal{F}^{j}_{\phi}$, where $\mathcal{F}^{j}_{R}$ and $\mathcal{F}^{j}_{\phi}$ satisfy the assumptions of Lemma 3.1, Lemma 3.2, respectively. For any direction $f\in\mathcal{F}:=\bigcup_{j\in\mathbb{Z}^{3}_{3}}\mathcal{F}^{j}$ let us define
@@
which is nothing but the the line in direction $f$ passing through the origin, but seen as subset of $\mathbb{T}^{3}$. Moreover, let $\eta$ be a positive constant and $\psi_{f}\in C^{\infty}(\mathbb{T}^{3},\mathbb{R})$ be a function such that:
@@ 3.2. Partition of unity and shifts
Next, let us fix $\nu\in(0,\frac{1-7\alpha}{4\alpha})$. Then, we introduce the parameter
(3.9) $\epsilon:=\begin{cases}C_{\epsilon}\lambda_{0}^{-1/2}\lambda_{1}^{-1/2},\,&q=0,\\ C_{\epsilon}\lambda_{q}^{-1/2}\lambda_{q+1}^{-1/2}\delta_{q}^{-1/4-\nu}\delta_{q+1}^{-1/4+\nu},\,&q\geq 1,\end{cases}\qquad C_{\epsilon}:=(C_{v}\overline{M}+L)^{-1}\min\left(\frac{1}{2},\frac{\inf_{j}N_{0}(\mathcal{F}_{R}^{j})}{6},\frac{\eta}{10\pi(L+1)}\right),$
which represents the size of the temporal partition, and
(3.10) $\mu^{-1}:=\begin{cases}\left\lceil\lambda_{0}^{1/2}\lambda_{1}^{1/2}\right\rceil,\,&q=0,\\ \left\lceil\lambda_{q}^{1/2}\lambda_{q+1}^{1/2}\delta_{q}^{1/4+\nu}\delta_{q+1}^{-1/4-\nu}\right\rceil,\,&q\geq 1,\end{cases}$
which represents instead the size of the spatial partition.
Given index sets
@@
(3.13) $\mathscr{F}[\mathbf{P}_{\leq 2^{j}}f](\xi):=\mathfrak{m}(\xi\,2^{-j})\mathcal{F}[f](\xi),\quad\mathscr{F}[\mathbf{P}_{>2^{j}}f](\xi):=[1-\mathfrak{m}(\xi\,2^{-j})]\mathscr{F}[f](\xi).$
Let us introduce the frequency cutoff parameter
[…]
@@
(3.16) $\mathbf{P}_{2^{j}}:=\mathbf{P}_{>2^{j-1}}-\mathbf{P}_{>2^{j}},$
that will be useful later. Then we define the spatial regularization of our velocity fields as:
@@
Since $\xi_{I}$, $I\in\mathcal{I}_{m}$ only depends on $m\in\mathbb{Z}$, we shall indifferently denote $\xi_{m}:=\xi_{I}$ hereafter.
@@ 3.3. Perturbation of the velocity
where the index $o$ stands for oscillatory. The amplitudes $a_{I}$ will be defined in the following subsection 3.4, and $\psi_{I}$ is given by
@@
The existence of such a family of shifts implies the following constraint on the parameters:
(3.21) $C_{\epsilon}\epsilon\,\mu\|\nabla u_{l}(t)\|_{0}\lesssim\frac{\eta}{\lambda_{q+1}}.$
Its proof is a simple adaptation of [DLK23, Proposition 3.5], with only minor modifications needed to take the randomness of $u_{l}$ into account, and is omitted for the sake of brevity. As in [DLK23], we can expand $\psi_{I},\psi^{2}_{I},\psi^{3}_{I}$ in Fourier series as follows:
(3.22) $\psi_{I}(x)=\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{b}_{I,k}e^{ik\cdot x},\quad\psi^{2}_{I}(x)=\mathring{c}_{I,0}+\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{c}_{I,k}e^{ik\cdot x},\quad\psi^{3}_{I}(x)=\mathring{d}_{I,0}+\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}\mathring{d}_{I,k}e^{ik\cdot x},$
@@
(3.23) $w_{o}(t,x)$ $=\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}b_{m,k}(t,x)e^{i\lambda_{q+1}\xi_{m}(t,x)\cdot k},$
(3.24) $(w_{o}\otimes w_{o})(t,x)$ $=c_{0}(t,x)+\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}c_{m,k}(t,x)e^{i\lambda_{q+1}\xi_{m}(t,x)\cdot k},$
(3.25) $\left(\frac{|w_{o}|^{2}}{2}w_{o}\right)(t,x)$ $=d_{0}(t,x)+\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}d_{m,k}(t,x)e^{i\lambda_{q+1}\xi_{m}(t,x)\cdot k},$
@@
(3.26) $b_{m,k}(t,x)=\sum_{I\in\mathcal{I}_{m}}\theta_{I}(t)\chi_{I}(\xi_{I}(t,x))a_{I}(t,x)\overline{f_{I}}(t,x)\,\mathring{b}_{I,k},$
(3.27) $c_{m,k}(t,x)=\sum_{I\in\mathcal{I}_{m}}\theta^{2}_{I}(t)\chi^{2}_{I}(\xi_{I}(t,x))a_{I}^{2}(t,x)\left(\overline{f_{I}}\otimes\overline{f_{I}}\right)(t,x)\,\mathring{c}_{I,k},$
(3.28) $d_{m,k}(t,x)=\sum_{I\in\mathcal{I}_{m}}\theta^{3}_{I}(t)\chi^{3}_{I}(\xi_{I}(t,x))a_{I}^{3}(t,x)\left(\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}\right)(t,x)\,\mathring{d}_{I,k},$
@@
(3.30) $c_{0}(t,x)=\sum_{I\in\mathcal{I}}\theta^{2}_{I}(t)\chi_{I}^{2}(\xi_{I}(t,x))a_{I}^{2}(t,x)\left(\overline{f_{I}}\otimes\overline{f_{I}}\right)(t,x)\fint_{\mathbb{T}^{3}}\psi_{I}(t,y)^{2}dy,$
(3.31) $d_{0}(t,x)=\sum_{I\in\mathcal{I}}\theta^{3}_{I}(t)\chi_{I}^{3}(\xi_{I}(t,x))a_{I}^{3}(t,x)\left(\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}\right)(t,x)\fint_{\mathbb{T}^{3}}\psi_{I}(t,y)^{3}dy.$
@@
(3.32) $w_{c}(t,x):=$ $\frac{1}{\lambda_{q+1}}\sum_{m\in\mathbb{Z}}\sum_{k\in\mathbb{Z}^{3}\setminus\{0\}}e_{m,k}(t,x)e^{i\lambda_{q+1}\xi_{m}(t,x)\cdot k},$
(3.33) $e_{m,k}(t,x)=$ $\sum_{I\in\mathcal{I}_{m}}\theta_{I}(t)\nabla\left(\chi_{I}(\xi_{I}(t,x))a_{I}(t,x)\right)\mathring{b}_{I,k}\times\left(\nabla\xi_{I}^{T}(t,x)\frac{ik\times f_{I}}{|k|^{2}}\right).$
@@ 3.4. Choice of the amplitudes
(3.34) $l_{\rm temp}:=\begin{cases}\lambda_{0}^{-1/2}\lambda_{1}^{-1/2},\,&q=0,\\ \lambda_{q}^{-1}\delta_{q}^{-2-\nu}\delta_{q+1}^{3/2+\nu},\,&q\geq 1,\end{cases}$
and let $\eta_{l_{\rm temp}}$ be the rescaled mollifier $\eta_{l_{\rm temp}}(t):=l^{-1}_{\rm temp}\eta(tl^{-1}_{\rm temp})$. The asymmetric choice for the support of $\eta$ guarantees that convolutions with $\eta_{l_{\rm temp}}$ preserve $\{\mathbb{F}_{t}\}_{t\geq 0}$ adaptedness. Then, define
(3.35) $R_{l}(t,x)$ $:=\int_{\mathbb{R}}(\mathbf{P}_{\leq l^{-1}}R_{q})\left(t+s,\mu(t;t+s,x)\right)\eta_{l_{\rm temp}}(s)ds,$
(3.36) $\phi_{l}(t,x)$ $:=\int_{\mathbb{R}}(\mathbf{P}_{\leq l^{-1}}\phi_{q})\left(t+s,\mu(t;t+s,x)\right)\eta_{l_{\rm temp}}(s)ds,$
@@
(3.39) $a_{I}(t,x):=|\overline{f_{I}}(t,x)|^{-2/3}2^{1/3}\Lambda_{f_{I}}^{1/3}\left(-\frac{\nabla\xi_{I}(t,x)\phi_{l}(t,x)}{\overline{M}^{3}\lambda_{q}^{-3\gamma/2}\delta_{q+1}^{3/2}}\right)\overline{M}(\lambda_{q}^{-3\gamma/2}\delta_{q+1}^{3/2})^{1/3}.$
Notice that $\epsilon\|u_{l}(t)\|_{1}<1/2$ implies $|\overline{f}_{I}(t,x)|>1/2$ by (3.29). Then, (3.7) and (3.31) plus the disjoint supports property gives
(3.40) $d_{0}$ $=\sum_{I\in\mathcal{I_{\phi}}}\theta^{3}_{I}\chi_{I}^{3}(\xi_{I})a_{I}^{3}\frac{|\overline{f_{I}}|^{2}}{2}\overline{f_{I}}$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t/\epsilon)\chi_{n}^{6}(\xi_{m})\nabla\xi_{m}^{-1}\sum_{\begin{subarray}{c}I\in\mathcal{I}_{\phi}\\ I=(m,n,f_{I})\end{subarray}}\overline{M}^{3}\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}\Lambda_{f_{I}}\left(-\frac{\nabla\xi_{I}\phi_{l}}{\overline{M}^{3}\lambda_{q}^{-\gamma}\delta_{q+1}^{3/2}}\right)f_{I}$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t/\epsilon)\chi_{n}^{6}(\xi_{m})(-\phi_{l})=-\phi_{l},$
where the second-to-last equality is due to Lemma 3.2 applied with $N_{0}=e$, while the last one is due to the space and time partitions of unity.
@@
(3.41) $a_{I}$ $:=\Gamma_{f}\left(\mathrm{Id}+(\nabla\xi_{I}\nabla\xi_{I}^{T}-\mathrm{Id})-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\rho^{1/2},$
@@
where $\Gamma_{f}$ and $N_{0}(\mathcal{F}_{R}^{j})$ are given by Lemma 3.1, $C>0$ is a constant depending on the $\Gamma_{f}$’s functions and $\sup_{f\in\mathcal{F}}|f|_{\infty}$, and
(3.43) $M_{m,n}(t,x):=\sum_{\begin{subarray}{c}J\in I_{\phi},\\ J=(m^{\prime},n^{\prime},f_{J}),\\ |m-m^{\prime}|_{\infty}\leq 1,\\ |n-n^{\prime}|_{\infty}\leq 1\end{subarray}}a_{J}^{2}(t,x)\chi_{J}^{2}(\xi_{J}(t,x))\theta_{J}^{2}(t)\left(\overline{f_{J}}\otimes\overline{f_{J}}\right)(t,x).$
@@
(3.44) $c_{0}$ $=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t/\epsilon)\chi_{n}^{6}(\xi_{m})\left[\sum_{I\in\mathcal{I}_{m,n}\cap\mathcal{I}_{R}}a_{I}^{2}\nabla\xi_{I}^{-1}(f_{I}\otimes f_{I})\nabla\xi_{I}^{-T}+M_{m,n}\right]$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t/\epsilon)\chi_{n}^{6}(\xi_{m})\left[\rho\nabla\xi_{I}^{-1}\left(\nabla\xi_{I}\nabla\xi_{I}^{T}-\frac{\nabla\xi_{I}(R_{l}-M_{m,n})\nabla\xi_{I}^{T}}{\rho}\right)\nabla\xi_{I}^{-T}+M_{m,n}\right]$
$=\sum_{(m,n)\in\mathbb{Z}\times\mathbb{Z}^{3}}\theta_{m}^{6}(t/\epsilon)\chi_{n}^{6}(\xi_{m})\left[\rho\mathrm{Id}-R_{l}\right]=\rho\mathrm{Id}-R_{l},$
where the second to last equality comes from Lemma 3.1 under the additional assumption